# Learning to Generate Novel Scientific Directions with Contextualized Literature-based Discovery

### 22 Jan 2024 v4

@@ 6 Conclusions and Future Directions
We propose a new contextualized literature-based discovery task for scientific hypothesis generation, where the goal is to generate texts suggesting new ideas grounded in rich natural language context for constraining the hypothesis search space. We collect and release a benchmark dataset by mining AI-related papers, and further annotate a corpus with human evaluations to test our approaches. We present a new framework for contextualized hypothesis generation by enriching generation models to retrieve inspirations from semantic similarity graphs, knowledge graphs, and citation networks.
Our experiments demonstrate that even with advanced LLMs, the task of generating natural language scientific hypotheses is highly challenging. Ideas predicted by our approaches tend to be incremental and with insufficient detail. Generating novel and meaningful scientific concepts and their compositions remains a fundamental problem Hope et al. (2023). Evaluation in this setting is also highly problematic, with a huge space of potentially plausible hypotheses formulated in natural language. One interesting direction is to expand C-LBD to incorporate a multimodal analysis of formulas, tables, and figures to provide a more comprehensive and enriched background context. We will also jointly train the retrieval and generation modules to improve retrieval performance.
@@ 7 Limitations
We discuss limitations extensively throughout the paper, namely in terms of evaluation challenges and data quality. Here we include additional details on limitations.
@@ 7.1 Limitations of Data Collection
We crawled papers with Semantic Scholar Academic Graph API from 1952 to June 2022. The number of available papers is limited by the data we crawled from Semantic Scholar Academic Graph. We also crawled papers from PubMed 1988 to 2023/08. We remove papers that are not English. We also remove papers where abstracts are not correctly parsed from paper PDFs. We will expand our models to papers written in other languages and other domains in the future.
@@ 7.2 Limitations of System Performance
Our dataset is based on state-of-the-art IE systems, which may be noisy. For instance, the coreference and SciSpacy abbreviation resolution models fail to link A2LCTC to Action-to-Language Connectionist Temporal Classification. The background context detection may also have errors: e.g., the sentence classification component fails to treat “For example, the language models are overall more positive towards the stock market, but there are significant differences in preferences between a pair of industry sectors, or even within a sector.” as background context. In our human-vetted gold data subset, we make sure to filter such cases, but they remain in the training data. SentenceBert Reimers and Gurevych (2019), and GPT3.5/GPT4 are not finetuned and might be biased towards pretraining datasets.
@@ 7.3 Limitations of Evaluation
The automatic metrics such as BLEU Papineni et al. (2002), ROUGE Lin (2004), BERTScore Zhang* et al. (2020), BARTScore Yuan et al. (2021), MRR, and Hits, might not be aligned with human judgment and be the best choices or our tasks. The gold subset and human evaluation are annotated by domain experts who might have biases in their selection.
Our setting uses a seed term taken from the ground truth as input, to emulate a scenario where a human provides guidance to an assistant model. Future work could explore methods in the setting without a seed term, an even harder task, or evaluate in an interactive setting with user-provided seed terms. In addition, while the seed is sampled from the ground truth, in our human-annotated gold subset we make sure that in no case the input context trivially leaks the output. This is also confirmed by low overall ROUGE-L scores.
@@ 7.4 Memorization Check
Carlini et al. (2023) shows that LLMs tend to memorize part of their training data, a well-known concern in evaluating current LLMs. Therefore, we examine the pretraining data of each model:
@@
  • GPT3.5: Based on the documentation666platform.openai.com/docs/model-index-for-researchers, GPT-3.5 series is pretrained on a combination of test and code from before Q4 2021.
  • GPT4: OpenAI (2023) shows that the GPT-4 series utilize most pertaining data before September 2021. Despite this, the pretraining and post-training data contain “a small amount” of more recent data. [^7] See footnote 10, page 10 of OpenAI (2023).
  • SciBERT: Beltagy et al. (2019) states that it is pretrained on papers published before 2019.
  • CS_RoBERTa: Gururangan et al. (2020) indicates that CS_RoBERTa is pretrained on papers papers before 2020.
Because we evaluate our models on papers published in 2022, the likelihood of test papers appearing in the pretraining corpora for the models is substantially reduced. We additionally performed a manual examination of GPT-4 memorization in our gold set based on 2022 ACL Anthology papers, by seeing if GPT-4 can complete information such as method names or generate text that strongly mimics the ground truth papers, and found no evidence of this occurring. The PMC-LLaMA Wu et al. (2023) is released in 2023/04 and our biochemical test set only includes PubMed papers after 2023/04.
@@ Appendix A Dataset Collection
We download ACL Anthology papers from 1952 to 2022 using Semantic Scholar Academic Graph API888www.semanticscholar.org/product/api. We filter out papers without abstracts and not written in English to obtain 67,408 papers. Our dataset has 58,874 papers before 2021, 5,946 papers from 2021, and 2,588 from 2022. We first use PL-Marker Ye et al. (2022) pretrained on SciERC Luan et al. (2018) to extract nodes belonging to six types: Task, Method, Evaluation Metric, Material, Other Scientific Terms, and Generic Terms. The model then predicts relations between nodes belonging to seven relation types: Used-for, Feature-of, Evaluate-for, Hyponym-of, Part-of, Compare, and Conjunction. Because we want to generate new ideas, we focus on used-for relations in computer science papers. Next, we use SciCo Cattan et al. (2021) with checkpoint from Hugging Face999huggingface.co/allenai/longformer-scico to obtain entity coreference to merge identical nodes. Then, we use ScispaCy Neumann et al. (2019) to perform unsupervised abbreviation detection to replace the abbreviation with a more informative long form. Finally, we perform scientific sentence classification Cohan et al. (2019)101010github.com/allenai/sequential_sentence_classification to classify sentences from the abstract into five categories including Background, Method, Objective, Other, and Result. We select sentences with labels of Background and Other as background context. During preprocessing, we only keep high-confidence outputs from IE models. Figure 3 shows an example from the non-canonicalized KG corpus constructed based on the IE systems pipeline. The statistics of the non-canonicalized KG are in Table 5.
@@
Type #Node #Relation
Train
70,462
61,606
Valid
11,970
8,576
Test
4,137
2,792
Table 5: Non-canonicalized KG statistics
@@
Appendix B Training and Finetuning details
@@ B.1 Inspiration Retrieval Module
The statistics of each neighbor type are in Table 7.
@@ B.2 Generation Module
Our T5/dual-encoder and their variants are built based on the Huggingface framework (Wolf et al., 2020)111111github.com/huggingface/transformers. We optimize those models by AdamW (Loshchilov and Hutter, 2019) with the linear warmup scheduler121212huggingface.co/docs/transformers/main_classes/optimizer_schedules#transformers.get_linear_schedule_with_warmup. Those models are finetuned on 4 NVIDIA A6000 48GB GPUs with distributed data parallel131313pytorch.org/tutorials/intermediate/ddp_tutorial.html. The training time for each model is about 10 hours.
B.2.1 T5 and its variants
We adopt the same input template for both tasks. For models without any neighbors, the input is the concatenation of the prompt $\mathcal{P}$ and background context $\mathcal{B}$ illustrated in §3.1 (i.e., $\mathcal{P}$ | context: $\mathcal{B}$). For models with neighbors, the input is $\mathcal{P}$ | retrieve: $n_{1},n_{2},..,n_{k}$ | context: $\mathcal{B}$, where $n_{1},...,n_{k}$ are retrieved neighbors. The input length is limited to $512$ tokens. For both tasks, we finetune our model based on T5-large with a learning rate of $6\times 10^{-6}$ and $\epsilon=1\times 10^{-6}$. The batch size is $8$ for each GPU. The maximum training epoch for all models is $10$ with $4$ patience.
Idea Sentence Generation
During decoding, we use beam-search to generate results with a beam size of 5 and a repetition penalty of $1.5$.
Idea Node Prediction
During decoding, because we want to compute HIT@10, we use beam-search to generate results with a beam size of $10$ and a number of beam groups of $10$. For the diversity beam search, we set the diversity penalty to $15$. For the constraint beam search, we constrain the output vocabulary to all entities available in the entire KG (i.e., the KG built on all papers in our corpus).
B.2.2 GPT and its variants
@@
We choose GPT3.5 davinci-003141414openai.com/api/ Brown et al. (2020) as our out-of-the-box causal language modeling baseline. We select $5$ instances from the training set as examples for the few-shot setting. We randomly select those examples for GPT3.5FS. For GPT3.5Retr, similar to semantic neighbors, we use all-mpnet-base-v2 from SentenceBert (Reimers and Gurevych, 2019), which performs best in semantic search to retrieve similar instances from the training set based on query $q$ in §3.1. The input length is limited to $2048$ tokens due to OpenAI API limits. We choose gpt-4-0314 as our GPT4 model. Our input for GPT4 is similar to GPT3.5.
Idea Sentence Generation
@@
For each selected example from the training set with forward relation, the template is “Consider the following context: $\mathcal{B}$ In that context, which $p$ can be used for $v$, and why? $s$, where $\mathcal{B}$ is the background context, $p$ is the target node type, $v$ is the seed term, and $s$ is the target idea sentence; for backward relation, the template is “Consider the following context: $\mathcal{B}$ In that context, which $p$ do we use $v$, and why? $s$. For selected examples with additional retrieval neighbors, we concatenate the following additional template to the $\mathcal{B}$: “The retrieval results are: $n_{1},n_{2},..,n_{k}$, where $n_{1},...,n_{k}$ are retrieved neighbors. For the final prompt, the template is similar to the above example template. However, the target idea sentence $s$ will not be included. We ask the model to generate $10$ outputs. We will select the best output and skip the empty output.
Idea Node Prediction
The template begins with a starting prompt. For the forward target relation, the template is “Suggest a $p$ that can be used for a natural language processing $v$; for the backward target relation, the template is “Suggest a $p$ that for a natural language processing $v$. For each selected example from the training set with forward relation, the template is “Consider the following context: $\mathcal{B}$ In that context, which $p$ can be used for $v$? $u$, where $u$ is the target node; for backward relation, the template is “Consider the following context: $\mathcal{B}$ In that context, which $p$ do we use $v$? $u$. For selected examples with additional retrieval neighbors, we also concatenate the following additional template to the $\mathcal{B}$: “The retrieval results are: $n_{1},n_{2},..,n_{k}$, where $n_{1},...,n_{k}$ are retrieved neighbors. For the final prompt, the template is similar to the above example template. However, the target node $u$ will not be included. We ask the model to generate $15$ outputs. We will select the top $10$ outputs and skip the empty output.
@@
Idea-Node Ground Truth
low-resource ner
[…]
Idea-Sentence Ground Truth
@@
Table 8: Example (from Zhou et al. (2022)) of retrieved neighbors for idea-sentence generation and idea-node generation. Neighbors similar to ground truth are underlined.
B.2.3 Dual-Encoder and its variants
[…]
$\mathcal{L}_{\mathrm{cl}}=\frac{e^{\left(\left(\phi(u,v)-\gamma\right)/\tau\right)}}{e^{\left(\left(\phi(u,v)-\gamma\right)/\tau\right)}+\sum_{v_{k}\in\mathcal{N}}e^{\left(\left(\phi(u,v_{k})\right)/\tau\right)}}$
where $\mathcal{N}$ is the negatives, $\tau$ is the temperature, and $\phi(u,v)=\cos(\mathbf{h}_{v},\mathbf{h}_{u})$.
B.3 In-context Contrastive Augmentation
B.3.1 T5 and its variants
Idea Sentence Generation
Idea Node Prediction
For this task, we randomly select $10$ scientific terms from scientific terms that appeared in the background context as in-context negatives. For example, in Figure 1, we treat scientific terms in the background context such as “pre-training”, “pre-training”, “PLMs”, etc.
B.3.2 Dual-Encoder and its variants
Similar to T5, the in-context negatives are scientific terms in the background context. We randomly select $5$ scientific terms that appeared in the background context as in-context negatives.
Model R-Lnormal-$\uparrow$ BARTnormal-$\uparrow$ BERTnormal-$\uparrow$
GPT3.5ZS
0.141
-6.123
0.630
GPT3.5FS
0.200
-5.690
0.651
GPT3.5Retr
0.202
-5.716
0.654
T5
0.260
-5.704
${}^{\dagger}$
@@ B.3.2 Dual-Encoder and its variants
0.262
-5.662
0.689
T5+SN+CL
0.264
-5.646
@@
0.258
-5.700
0.687
T5+CT+CL $*{}^{*}$$*{}^{*}$ ${}^{\dagger}$
Table 9: Automatic evaluation results on idea sentence generation (full dataset). BART denotes BARTscore. $*{}^{*}$ denotes that the observed differences are statistically significant ($p\leq 0.05$).
3FS 3Rt 3FS+CT 3FS+KG 4 4+KG T5 T5+SN
Len
24.3
24.1
23.1
29.7
58.9
58.9
19.6
22.5
Table 10: Average length of each system output receives for human evaluation set. “Len” denotes the average length. “3FS” refers to the GPT3.5FS. “3Rt” refers to the GPT3.5Retr. “4” refers to the GPT4FS. “T5+SN” refers to the T5+SN+CL.
Appendix C Evaluation Metrics
We use the official implementation of BARTScore (Yuan et al., 2021) and its corresponding checkpoint 181818github.com/neulab/BARTScore for both tasks. We use BERTScore Zhang* et al. (2020) with SciBERT checkpoint for both tasks. The hash of the checkpoint is allenai/scibert_scivocab_uncased_L8 _no-idf_version=0.3.12(hug_trans=4.19.2). The automatic evaluation on idea sentence generation for the full dataset without any filters is in Table 9
Appendix D Length for idea-sentence generation
We computed the length of the idea-sentence generation for each baseline within the human evaluation set, with results provided in Table 10. We find that the results from GPT-4 are more than twice as long as those from other baselines, suggesting GPT4 outputs contain more information.
Appendix E Automated evaluation for idea-node prediction
Model MRRnormal-$\uparrow$ HIT@1normal-$\uparrow$ HIT@5normal-$\uparrow$ HIT@10normal-$\uparrow$
GPT3.5FS
0.007
0.004
0.010
0.014
GPT3.5Retr
0.003
0.000
0.000
0.027
SciBERT
@@ Appendix E Automated evaluation for idea-node prediction
0.108
0.138
T5+Div+CL
0.029
0.019
0.041
0.054
[…]
GPT3.5FS+SN
0.007
0.004
0.011
0.013
SciBERT+SN
0.080
0.046
0.110
0.143
CSRoBERTa+SN
0.073
0.041
0.099
0.132
T5+Div+SN
0.029
0.019
0.042
0.052
[…]
GPT3.5FS+KG
0.007
0.004
0.011
0.013
SciBERT+KG
0.079
0.044
0.111
0.144
CSRoBERTa+KG
0.071
0.040
0.101
0.130
T5+Div+KG
0.029
0.020
0.041
0.053
[…]
GPT3.5FS+CT
0.007
0.003
0.013
0.017
SciBERT+CT ${}^{\dagger}$

### 17 Feb 2024 v5

@@ 5 Conclusions and Future Directions
We propose a new setting, model and comprehensive evaluation for scientific hypothesis generation with language models that are grounded in literature and optimized for novelty. We present a new framework named SciMON in which models take background problem contexts and provide suggestions that are novel while based on literature. Models retrieve inspirations from semantic similarity graphs, knowledge graphs, and citation networks. We introduce a new iterative novelty boosting mechanism that helps large language models (LLMs) such as GPT-4 generate more novel ideas by explicitly comparing ideas to prior work and refining them. Our experiments demonstrate that the task of generating natural language scientific hypotheses is highly challenging. While our methods improve upon baseline LLMs, generated ideas tend to be incremental and with insufficient detail. Generating novel and meaningful scientific concepts and their compositions remains a fundamental problem Hope et al. (2023). Evaluation in this setting is also highly challenging, with a huge space of potentially plausible hypotheses formulated in natural language. One interesting direction is to expand SciMON with a multimodal analysis of formulas, tables, and figures to provide a more comprehensive background context.
@@ 6 Limitations
We discuss limitations extensively throughout the paper, such as in terms of evaluation challenges and data quality. Here we include additional details on limitations.
@@ 6.1 Limitations of Data Collection
We crawled papers with Semantic Scholar Academic Graph API from 1952 to June 2022. The number of available papers is limited by the data we crawled from the Semantic Scholar Academic Graph. We also crawled papers from PubMed 1988 to 2024/01. We remove papers that are not English. We also remove papers where abstracts are not correctly parsed from paper PDFs. We will expand our models to papers written in other languages and other domains in the future.
@@ 6.2 Limitations of System Performance
Our dataset is based on state-of-the-art IE systems, which may be noisy. For instance, the coreference and SciSpacy abbreviation resolution models fail to link A2LCTC to Action-to-Language Connectionist Temporal Classification. The background context detection may also have errors: e.g., the sentence classification component fails to treat “For example, the language models are overall more positive towards the stock market, but there are significant differences in preferences between a pair of industry sectors, or even within a sector.” as background context. In our human-vetted gold data subset, we make sure to filter such cases, but they remain in the training data. SentenceBert Reimers and Gurevych (2019), and GPT3.5/4 are not finetuned and might be biased towards pretraining datasets. The idea novelty boosting method is limited by the quality of retrieval models. Better retrieval models may be explored in the future. Due to hardware constraints, we mainly investigated models with up to 7 billion parameters. Due to API change and model randomness, our GPT3.5/4 results might not be easily reproducible.
@@ 6.3 Limitations of Evaluation
We recruit annotators from Ph.D. students; their opinions may differ from annotators who have different levels of domain knowledge. Our setting uses a seed term taken from the ground truth as input, to emulate a scenario where a human provides guidance to an assistant model. Future work could explore methods in the setting without a seed term, an even harder task, or evaluate in an interactive setting with user-provided seed terms. In addition, while the seed is sampled from the ground truth, in our human-annotated gold subset, we make sure that in no case does the input context trivially leak the output.
@@ 6.4 Memorization Check
Carlini et al. (2023) reports that LLMs tend to memorize part of their training data, a well-known concern in evaluating current LLMs. Therefore, we examine the pretraining data of each model:
@@
  • GPT3.5: Based on the documentation,999platform.openai.com/docs/model-index-for-researchers GPT-3.5 series is pretrained on a combination of test and code from before Q4 2021.
  • GPT4: OpenAI (2023) shows that the GPT-4 checkpoint we used utilizes most pertaining data before September 2021. Despite this, the pretraining and post-training data contain “a small amount” of more recent data. [^10] See footnote 10, page 10 of OpenAI (2023).
Because we evaluate our models on papers published in 2022, the likelihood of test papers appearing in the pretraining corpora for the models is substantially reduced. We additionally performed a manual examination of GPT-4 memorization in our gold set based on 2022 ACL Anthology papers, by seeing if GPT-4 could complete information such as method names or generate text that strongly mimics the ground truth papers, and found no evidence of this occurring. The Meditron-7b Chen et al. (2023) uses PubMed with a cut-off in August 2023, and our biochemical test set only includes PubMed papers after 2023/08.
@@ Appendix A Dataset Collection
A.1 NLP Dataset Collection
We download ACL Anthology papers from 1952 to 2022 using Semantic Scholar Academic Graph API.111111www.semanticscholar.org/product/api We filter out papers without abstracts and not written in English to obtain 67,408 papers. Our dataset has 58,874 papers before 2021, 5,946 papers from 2021, and 2,588 from 2022. We first use PL-Marker Ye et al. (2022) pretrained on SciERC Luan et al. (2018) to extract nodes belonging to six types: Task, Method, Evaluation Metric, Material, Other Scientific Terms, and Generic Terms. The model then predicts relations between nodes belonging to seven relation types: Used-for, Feature-of, Evaluate-for, Hyponym-of, Part-of, Compare, and Conjunction. Because we want to generate new ideas, we focus on used-for relations in papers. Next, we use SciCo Cattan et al. (2021) with checkpoint from Hugging Face121212huggingface.co/allenai/longformer-scico to obtain entity coreference to merge identical nodes. Then, we use ScispaCy Neumann et al. (2019) to perform unsupervised abbreviation detection to replace the abbreviation with a more informative long form. Finally, we perform scientific sentence classification Cohan et al. (2019)131313github.com/allenai/sequential_sentence_classification to classify sentences from the abstract into five categories including Background, Method, Objective, Other, and Result. We select sentences with labels of Background and Other as background context. During preprocessing, we only keep high-confidence outputs from IE models. Figure 4 shows an example of the IE systems pipeline.
@@ A.1 NLP Dataset Collection
A.2 Biochemical Dataset Collection
We collect PubMed papers from 1988 to 2024 using Entrez Programming Utilities API141414www.ncbi.nlm.nih.gov/books/NBK25501/ for the following topics, including Yarrowia, Saccharomyces cerevisiae, Issatchenkia orientalis, and Rhodosporidium toruloides. We use PubTator 3 Islamaj et al. (2021); Wei et al. (2022); Luo et al. (2023); Wei et al. (2023); Lai et al. (2023). The PubTator 3 performs named entity recognition, relation extraction, entity coreference and linking, and entity normalization for the abstracts in the dataset. PubTator 3 identifies bio entities belonging to seven types: gene, chemical, chromosome, cell line, variant, disease, and speciesl and relations belonging to 13 types: associate, cause, compare, convert, contract, drug interact, inhibit, interact, negative correlate, positive correlate, prevent, stimulate, and treat. Finally, we use a sentence classifier trained on CODA-19 Huang et al. (2020) to classify sentences in abstracts into background, purpose, method, finding, and other. We select sentences with labels of background as background context and remove sentences with labels of other. We treat the rest sentences that have at least one entity as the target sentence. We only keep samples with low similarity between background context and corresponding ground truth sentences. [^15] The similarity is calculated with all-mpnet-base-v2. Our final dataset has 4,767 papers before 2023/02, 642 papers from 2023/02 to 2023/08, and 299 papers after 2023/08.
Appendix B Finetuning and Automated Evaluation details
@@ B.1 Inspiration Retrieval Module
The statistics of each inspiration type are in Table 7. Table 8 shows sample retrieved inspirations.
@@ B.2 Generation Module
Our T5 model and their variants are built based on the Huggingface framework (Wolf et al., 2020).161616github.com/huggingface/transformers We optimize those models by AdamW (Loshchilov and Hutter, 2019) with the linear warmup scheduler.171717huggingface.co/docs/transformers/main_classes/optimizer_schedules#transformers.get_linear_schedule_with_warmup Those models are finetuned on 4 NVIDIA A6000 48GB GPUs with distributed data parallel.181818pytorch.org/tutorials/intermediate/ddp_tutorial.html The training time for each model is about 10 hours.
B.2.1 In-Context Learning
@@
We choose GPT3.5 davinci-003191919openai.com/api/ Brown et al. (2020) as our out-of-the-box causal language modeling baseline. We select $5$ instances from the training set as examples for the few-shot setting. We randomly select those examples for GPT3.5FS. For GPT3.5Retr, similar to semantic neighbors, we use all-mpnet-base-v2 from SentenceBert (Reimers and Gurevych, 2019), which performs best in semantic search to retrieve similar instances from the training set based on query $q$ in §3.1. The input length is limited to $2048$ tokens due to OpenAI API limits. We choose gpt-4-0314 as our GPT4 model. Our input for GPT4 is similar to GPT3.5.
@@
For each selected example from the training set with forward relation, the template is “Consider the following context: $\mathcal{M}$ In that context, which $p$ can be used for $v$, and why? $\mathcal{T}$, where $\mathcal{M}$ is the background context, $p$ is the target node type, $v$ is the seed term, and $\mathcal{T}$ is the target idea sentence; for backward relation, the template is “Consider the following context: $\mathcal{M}$ In that context, which $p$ do we use $v$, and why? $s$. For selected examples with additional retrieval inspirations, we concatenate the following additional template to the $\mathcal{M}$: “The retrieval results are: $i_{1},\ldots,i_{k}$, where $i_{1},\ldots,i_{k}$ are retrieved inspirations. For the final prompt, the template is similar to the above example template. However, the target sentence $\mathcal{T}$ will not be included. We ask the model to generate $10$ outputs. We will select the best output and skip the empty output.
B.2.2 Fine Tuning
Given input without any inspirations, the input combines the prompt $\mathcal{P}$ and context $\mathcal{M}$ as shown in §3.1 (i.e., $\mathcal{P}$ | context: $\mathcal{M}$). Given input with inspirations, the input is $\mathcal{P}$ | retrieve: $i_{1},\ldots,i_{k}$ | context: $\mathcal{M}$, with $i_{1},\ldots,i_{k}$ as retrieved inspirations. The input length is limited to $512$ tokens. For both tasks, we finetune our model based on T5-large with a learning rate of $6\times 10^{-6}$ and $\epsilon=1\times 10^{-6}$. The batch size is $8$ for each GPU. The maximum training epoch for all models is $10$ with $4$ patience. During decoding, we use beam-search to generate results with a beam size of 5 and a repetition penalty of $1.5$.
In-context Contrastive Augmentation
B.2.3 Biochemical Case Study
Our Meditron-7b Chen et al. (2023) and its variants are built based on the Huggingface framework (Wolf et al., 2020).202020github.com/huggingface/transformers We use its epfl-llm/meditron-7b as the base model. We finetune those models with a learning rate of $2\times 10^{-6}$ and $\epsilon=5\times 10^{-8}$. The maximum training epoch for all models is $5$. All models are fine-tuned on 4 NVIDIA A100 80 GB GPUs with Fully Sharded Data Parallel. [^21] https://huggingface.co/docs/accelerate/usage_guides/fsdp The training time for each model is about 20 hours.
@@ B.2.3 Biochemical Case Study
Ground Truth
@@
Table 8: Example (from Zhou et al. (2022)) of retrieved inspirations. Inspirations similar to ground truth are underlined.
B.3 Automated Evaluation
We use BERTScore Zhang* et al. (2020) with SciBERT checkpoint for both tasks. The hash of the checkpoint is allenai/scibert_scivocab_uncased_L8 _no-idf_version=0.3.12(hug_trans=4.19.2). The automated evaluation results are in Table 9.
Figure 5: Gold subset annotation interface
Subset Challenging Gold
@@ B.3 Automated Evaluation
${}^{\dagger}$
${}^{\dagger}$
${}^{\dagger}$
${}^{\dagger}$
T5+SN+CL ${}^{\dagger}$
${}^{\dagger}$
${}^{\dagger}$